A large rectangular prism has a surface area of 576 square feet. The dimensions of a small rectangular prism are 1/3 the size of the dimensions of the large rectangular prism. What is the surface area of the small rectangular prism?

Answer by fractalier(6550)

The surface area is proportional to the square of the ratio of the sides. So if the smaller one is 1/3 the size on each side, the surface area is (1/3)^2 = 1/9 as great.
Thus, 1/9 of 576 is 576/9 = 64 square feet